OpenStudy (anonymous):

Simplify: w/6^2 ?? can someone explain this to me please?

OpenStudy (anonymous):

What is the title of the chapter or whatever?

OpenStudy (solomonzelman):

what do you want to do with w/6^2 ?

OpenStudy (anonymous):

Its properties of exponents and i just need to figure out the answer i guess? im not very good at math

OpenStudy (solomonzelman):

is it (w/6)^2 ?

OpenStudy (anonymous):

yes it is

OpenStudy (solomonzelman):

so, \(\large\color{black}{ \displaystyle \left(\frac{ A}{B}\right)^n=\frac{A^n}{B^n} }\)

OpenStudy (solomonzelman):

use that, but with \(\large\color{black}{ \displaystyle \left(\frac{w}{6}\right)^2 }\)

OpenStudy (solomonzelman):

your A is w your B is 6 your n is 2

OpenStudy (anonymous):

\[\frac{ w^2 }{ 6^2 }\]

OpenStudy (anonymous):

Would the answer be w^2/36?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh thank you guys for helping

OpenStudy (solomonzelman):

yes, correct